# Rational motives on pro-algebraic stacks

### v3

@@ Abstract
We define an $\infty$-category of rational motives for cofiltered limits of algebraic stacks, so-called pro-algebraic stacks. We show that it admits a $6$-functor formalism for certain classes of morphisms. Our main example is the stack of displays, closely related to the stack of Barsotti-Tate groups, seen as a pro-algebraic stack via its truncations. On the way we deal with base change along non-finite type morphisms such as non-finite Galois extensions.
@@ 1 Introduction
The study of algebraic cycles is important in algebraic geometry, number theory and representation theory. Over the years there have been many conjectures and results concerning algebraic cycles, that we could not list in its detail in this article. Not only have cycles on varieties been studied but also the study of equivariant cycles has become important, for example for the construction of higher Gross-Zagier formulas by Yun and Zhang (cf. [YZ17]).
For the study of equivariant cycles, Edidin and Graham constructed equivariant Chow groups (cf. [EG98]) following the idea of Totaro (cf. [Tot99]). In his PhD-thesis Brokemper gave a computation of the Chow ring of the stack of truncated Barsotti-Tate groups ([Bro18]). The main idea was to relate the Chow groups of the stack of truncated Barsotti-Tate groups to the stack of so-called truncated displays due to Lau ([Lau13]). Let us denote the stack of $n$-truncated displays by $\textup{Disp}_{n}$. This is an Artin stack over $\mathbb{F}_{p}$. For each $n\geq 1$, we get a map $\tau_{n}\colon\textup{Disp}_{n+1}\rightarrow\textup{Disp}_{n}$. Brokemper showed that this map admits a factorization
[…]
where $\alpha$ is an affine bundle and $\beta$ is a map of quotient stacks $\left[X/G\right]\rightarrow\left[X/H\right]$ induced by an extension of a unipotent group, i.e. an exact sequence
$0\rightarrow U\rightarrow G\rightarrow H\rightarrow 0,$
where $U$ is unipotent. By homotopy invariance, we therefore have $A^{*}(\textup{Disp}_{n},*)\cong A^{*}(\textup{Disp}_{1},*)$. One can show that $\textup{Disp}_{1}$ has a particular representation as a quotient stack, which allows Brokemper to compute its Chow groups explicitly. An interesting application of this result is the following. Let $\textup{BT}_{n}$ denote the stack of $n$-truncated Barsotti-Tate groups. Then there exists a smooth morphism $\phi_{n}\colon\textup{BT}_{n}\rightarrow\textup{Disp}_{n}$, that is compatible with truncations and is an equivalence on geometric points (cf. [Lau13]). This morphism induces an isomorphism on the level of rational Chow groups and an injection on the level of integral Chow groups (cf. [Bro18]). Rationally, this is clear as $\phi_{n}$ is a universal homeomorphism and thus induces an equivalence of motives. Thus, for any $n\in\mathbb{N}$, we have
$A^{*}(\textup{BT}_{n},*)_{\mathbb{Q}}\cong A^{*}(\textup{Disp}_{1},*)_{\mathbb{Q}}.$
Due to the work of Levine, we know that the Chow groups can be realized by motivic cohomology (cf. [Lev]). Rationally the same is true, even for equivariant Chow groups but let us be a bit more precise. For this paragraph, we assume that every Artin stack, and hence also every scheme, is of finite type over a field. The category of Beilinson motives satisfies étale descent (cf. [CD19]). Using this, we can follow the construction of Richarz and Scholbach to define $\textup{DM}(X)$, the $\infty$-category of rational motives on an Artin stack $X$ with smooth atlas $\tilde{X}$, via the formula
[…]
@@
(cf. [RS20]). Equivalently, as shown in op. cit., we can define $\textup{DM}(X)$ via right Kan extension along the Yoneda embedding. This approach shows how to define $\textup{DM}(F)$ for any prestack $SSF\colon\textup{(Ring)}\rightarrow\SS$ (here $SS\SS$ denotes the $\infty$-category of Kan-complexes).
One particular question, that is the initial motivation for this article, is what happens with the motivic cohomology of the non-truncated stack of displays, i.e. the motivic cohomology of $\textup{Disp}\coloneqq\lim_{n}\textup{Disp}_{n}$. The motivic cohomology of Disp should be equal to the motivic cohomology of $\textup{Disp}_{1}$ since we enlarge $\textup{Disp}_{1}$ in a way that cannot be seen in $\mathbb{A}^{1}$-homotopy theory. However the computation of motivic cohomology via right Kan extension seems to more complicated. Instead of this approach, we will give an alternative definition of DM for cofiltered limits of algebraic stacks of finite type, so-called pro-algebraic stacks.
So let us fix a filtered category ${\mathcal{I}}$ and for the sake of this introduction let us fix a field $k$ and $S\coloneqq\operatorname{Spec}(k)$.
@@
A $-{\mathcal{I}}$-algebraic stack is a map $x\colon{\mathcal{I}}^{\operatorname{op}}\rightarrow\textup{PreStk}_{S}$, i.e. a cofiltered systems of prestacks, such that each $X_{i}\coloneqq x(i)$ is an Artin stack locally of finite type over $S$. We denote such stacks as $(x,X)$, where $X\coloneqq\lim_{{\mathcal{I}}}x$. The transition maps are denoted by $x_{ij}\colon X_{j}\rightarrow X_{i}$ for any map $i\rightarrow j\in{\mathcal{I}}$ (this is an abusive notation but for convenience the ambiguity can be neglected). We say that $(x,X)$ is
  1. 1.
    classical if $x_{ij}$ is affine for all $i\rightarrow j\in{\mathcal{I}}$,
  2. 2.
    monoidal if $x_{ij}$ is smooth for all $i\rightarrow j\in{\mathcal{I}}$, and
  3. 3.
A morphism $f\colon(x,X)\rightarrow(y,Y)$ of $-{\mathcal{I}}$-algebraic stacks is a morphism of diagrams. We denote by $f_{i}\colon X_{i}\rightarrow Y_{i}$ the induced map. We call a morphism $f\colon(x,X)\rightarrow(y,Y)$ of $-{\mathcal{I}}$-algebraic stacks cartesian if each square
[…]
is a pullback diagram. We call a morphism $f\colon(x,X)\rightarrow(y,Y)$ of $-{\mathcal{I}}$-algebraic stacks adjointable if the morphism $y_{ij}^{*}f_{i*}\rightarrow f_{j*}x_{ij}^{*}$ induced by each square above is an equivalence.
Let us mention $2$ examples of pro-algebraic stacks.
Example 2.
  1. (1)
    Let $k$ be a field and $L/k$ be a (not necessarily finite) Galois extension. Then the system of finite Galois extensions ${\mathcal{G}}_{k}^{L}$ of $k$ containing $L$ induces a classical monoidal pro-${\mathcal{G}}_{k}^{L}$-algebraic stack. Similarly, via base change, we can define for any algebraic $k$-stack $f\colon{\mathfrak{X}}\rightarrow\operatorname{Spec}(k)$ a pro-${\mathcal{G}}_{k}^{L}$-algebraic stack $(q_{{\mathfrak{X}}},{\mathfrak{X}}_{L})$ that comes with a cartesian map to $(q,\operatorname{Spec}(L))$.
  2. (2)
    The $-\mathbb{N}$-algebraic stack $(\tau,\textup{Disp})$ is by the discussion above strict and monoidal but not classical.
Example (1) above will show that our formalism of motives on pro-algebraic stacks allows us to work with non-finite type morphisms of schemes and get for example a smooth base change in this setting. Example (2) was discussed in the beginning of this introduction and will show how to define motives for stacks that are not necessarily algebraic.
Let us also give $2$ interesting examples of adjointable morphisms.
Example 3.
  1. (1)
    Let $f\colon(x,X)\rightarrow(y,Y)$ be a cartesian morphism of monoidal $-{\mathcal{I}}$-algebraic stacks, then by smooth base change $f$ is adjointable.
  2. (2)
    Let $f\colon(x,X)\rightarrow(y,Y)$ be a morphism of $-{\mathcal{I}}$-algebraic stacks. Assume that each $y_{ij*}$ is fully faithful and $(x,X)$ is strict, then simple computation shows that $f$ is adjointable. An example of such a situation is the structure map $(\tau,\textup{Disp})\rightarrow\operatorname{Spec}(\mathbb{F}_{p})$, where $(\tau,\textup{Disp})$ is as in Example 2.
@@
We define the $\infty$-category of rational motives on a $-{\mathcal{I}}$-algebraic stack $(x,X)$ to be
$\textup{DM}(x,X)\coloneqq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{*}}\textup{DM}(X_{i}).$
Note that if $(x,X)$ is classical, then $X$ is in fact representable by an algebraic stack and we have $\textup{DM}(x,X)\simeq\textup{DM}(X)$ (see Proposition 3.1). Also, with this definition, it is clear that $\textup{DM}(x,X)$ admits a closed symmetric monoidal structure and is compactly generated. One of the main motivations behind our choice of definition is that we obtain a (motivic) $6$-functor formalism for certain classes of morphisms.
Theorem 5.
Let $f\colon(x,X)\rightarrow(y,Y)$ be a morphism of monoidal $-{\mathcal{I}}$-algebraic stacks. Then there is an adjunction
${f^{*}\colon\textup{DM}(y,Y)}$${\textup{DM}(x,X)\colon f_{*}}$
functorial in $(x,X)$ and $(y,Y)$, where $f^{*}$ is symmetric monoidal and preserves compact objects. Moreover, the functor DM together with $f^{*}$ defines a pullback formalism, i.e. if $f$ is smooth then $f^{*}$ admits a left adjoint $f_{\sharp}$ that satisfies base change and projection formula.
Assume now that ${\mathcal{I}}\simeq\mathbb{N}_{0}$ and $f$ is adjointable or that ${\mathcal{I}}$ is arbitrary and $f$ is cartesian. Then we can upgrade $f^{*}$ to be part of a $6$-functor formalism in the following sense. There exists a functorial adjunction
${f_{!}\colon\textup{DM}(x,X)}$${\textup{DM}(y,Y)\colon f^{!}.}$
Further, the adjunctions above satisfy the following properties.
  1. (F1)
    If $f$ is proper, we have a canonical equivalence $f_{*}\simeq f_{!}$.
  2. (F2)
    The functors $f_{!}$ and $f^{!}$ commute with arbitrary colimits.
  3. (F3)
    (Base change) For any pullback square
    ${(w,W)}$${(z,Z)}$${(x,X)}$${(y,Y)}$$\scriptstyle{f^{\prime}}$$\scriptstyle{g^{\prime}}$$\scriptstyle{g}$$\scriptstyle{f}$
    of monoidal pro-algebraic stacks with morphisms satisfying (i) or (ii), the exchange morphisms
    $f^{*}g_{!}\rightarrow g^{\prime}_{!}f^{\prime*}$
    $g^{!}f_{*}\rightarrow f^{\prime}_{*}g^{\prime!}$
    are equivalences.
  4. (F4)
  5. (M1)
    (Homotopy invariance) Let us define $(q,\mathbb{A}^{1}_{X})$ via the system $n\mapsto\mathbb{A}^{1}_{X_{i}}$. Then $p^{*}$ is fully faithful, where $p\colon(q,\mathbb{A}^{1}_{X})\rightarrow(x,X)$ denotes the projection.
  6. (M2)
  7. (M3)
    (Purity) We have a canonical equivalence $f^{!}\simeq\operatorname{\textup{Th}}(\Omega_{f_{0}})1_{X}\otimes f^{*}$.
  8. (M4)
  9. (M5)
    $i_{!}i^{!}\rightarrow\operatorname{id}\rightarrow j_{*}j^{*}$
    $j_{!}j^{!}\rightarrow\operatorname{id}\rightarrow i_{*}i^{*}.$
Even though the proof is formal, this theorem allows us to explicitly compute the motive of (strict) $-{\mathcal{I}}$-algebraic stacks and their motivic cohomology. For example, if $(x,X)$ is a strict monoidal $-{\mathcal{I}}$-algebraic stack with a map $f\colon(x,X)\rightarrow Y$, where $Y$ is an algebraic stack seen as a constant diagram, then
$f_{!}f^{!}1_{Y}\simeq f_{0!}f^{!}_{0}1_{Y},$
as expected if ${\mathcal{I}}$ has an initial object 00.
Remark 6.
Let us quickly mention that we develop the theory of rational motives not only for pro-algebraic stacks. Theorem 5 works similarly for any (motivic) six functor formalisms that are compactly generated if we restrict us to $-{\mathcal{I}}$-schemes (analogously defined to the stack case). If further the (motivic) six functor formalism satisfies étale descent, then we this also holds for $-{\mathcal{I}}$-algebraic stacks.
Let us define absolute motivic cohomology as follows.
Definition 7.
Let $(x,X)$ be a monoidal $-{\mathcal{I}}$-algebraic stack and $n,m\in\mathbb{Z}$. Then we define the absolute motivic cohomology of $(x,X)$ in degree $(n,m)$ by
$H^{n,m}(X,\mathbb{Q})\coloneqq\operatorname{Hom}_{\textup{DM}(x,X)}(1_{(x,X)},1_{(x,X)}(n)[m]),$
where we define Tate-twists according to Theorem 5.
We will see in Proposition 3.10 that if $(x,X)$ is classical and $X$ is represented by an algebraic stack locally of finite type over $S$, then $H^{n,m}(X,\mathbb{Q})\cong\operatorname{Hom}_{\textup{DM}(X)}(1_{X},1_{X}(n)[m])$ recovering the usual definition. Even in the non-classical case, following Brokemper’s thesis (cf. [Bro16]), we will see in Example 3.12 that if $k$ is a field, $L$ and algebraic extension and $Z$ a $k$-scheme, then $H^{n,m}(Z_{L},\mathbb{Q})\cong A^{n}(Z_{L},2n-m)_{\mathbb{Q}}$ as expected.
Let us now come back to the setting of the beginning of this introduction, i.e. Example 2 (2). Recall, that the truncation $\tau_{n}\colon\textup{Disp}_{n+1}\rightarrow\textup{Disp}_{n}$ is a composition of an affine bundle and a map of quotient stacks induced by an extension of a unipotent group. Then the following observation follows immediately from construction and the author’s computations of Tate motives on quotients by Frobenius conjugation (cf. [Yay23a])
@@
Let $f\colon\textup{Disp}\rightarrow\operatorname{Spec}(\mathbb{F}_{p})$ denote the structure map. Then $f_{*}f^{*}1_{\mathbb{F}_{p}}\simeq f_{1*}f_{1}^{*}1_{\mathbb{F}_{p}}$ and $f_{\sharp}f^{*}1_{\mathbb{F}_{p}}\simeq f_{1\sharp}f_{1}^{*}1_{\mathbb{F}_{p}}$ are contained in the ind-completion of the full stable subcategory of $\textup{DM}(S)$ generated by Tate-motives.
The equivalence $f_{*}f^{*}1_{\mathbb{F}_{p}}\simeq f_{1*}f_{1}^{*}1_{\mathbb{F}_{p}}$ stated in the above theorem also allows us to compute the motivic cohomology of Disp after truncation (this has been done by Brokemper, cf. [Bro18]).
@@
We start by defining pro-algebraic stacks and listing some examples. We define the $\infty$-category of rational motives on pro-algebraic stacks.
Afterward, we prove our main theorem, that on certain classes of morphisms, this definition yields a (motivic) $6$-functor formalism. We do this in a very general setting considering pullback formalisms, coefficient systems and abstract $6$-functor formalisms.
Lastly, we apply our theory to the three examples mentioned in the introduction.

### v4

# Motives on pro-algebraic stacks

@@ Abstract
We define an $\infty$-category of rational motives for inverse limits of algebraic stacks, so-called pro-algebraic stacks. We show that it admits a $6$-functor formalism for certain classes of morphisms. On pro-schemes, we show that this 6-functor formalism is in the sense of Liu-Zheng. This theory yields an approach to the theory of motives for non-representable algebraic stacks and non-finite type morphisms.
@@ 1 Introduction
Motivation: Barsotti–Tate groups have been central to many studies in arithmetic geometry. Consider the stack BT of Barsotti–Tate groups over $\operatorname{Spec}(\mathbb{F}_{p})$. Geometrically, BT is difficult to understand, as it is not an algebraic stack. However, we can express BT as a pro-system of algebraic stacks, $\textup{BT}\simeq\lim_{n\geq 0}\textup{BT}_{n}$, where $\textup{BT}_{n}$ denotes the algebraic stack of $n$-truncated Barsotti–Tate groups. Given the importance of BT, it is natural to ask about its “motive” in $\textup{DM}(\mathbb{F}_{p},\mathbb{Q})$. Since BT is not algebraic, this question does not a priori make sense, though one can formally define
$M(\textup{BT})\coloneqq\mathop{\textup{colim}}_{n\geq 0}M(\textup{BT}_{n}).$
Alternatively, we could formally extend $\textup{DM}^{!}_{\mathbb{Q}}$ to arbitrary prestacks and define $M(\textup{BT})$ as the !!-pull/push of the unit. However, there is no general reason why these two definitions should agree. This issue is often referred to as continuity. If $S=\lim_{i}S_{i}$ is a projective system of noetherian finite-dimensional schemes with affine transition maps, then
$M(S)\cong\mathop{\textup{colim}}_{i}M(S_{i})$
provided that $S$ is also noetherian of finite dimension [CD19, Thm. 14.3.1].
Returning to BT, both of the above definitions seem reasonable. However, the formal extension $\textup{DM}^{!}_{\mathbb{Q}}$ does not come with the six-functor formalism one would typically expect in a theory of motives. For computations, such a formalism would be very useful. This becomes clearer when considering the stack Disp of so-called displays [Lau13]. This stack also admits a presentation $\textup{Disp}\simeq\lim_{n\geq 0}\textup{Disp}_{n}$ as a pro-system of algebraic stacks over $\operatorname{Spec}(\mathbb{F}_{p})$, and we have a map of pro-systems of algebraic stacks $\phi\colon\textup{BT}\to\textup{Disp}$ [Lau13]. This morphism is level-wise smooth and a universal homeomorphism. The motive of each $\textup{Disp}_{n}$ is in fact Tate (cf. Theorem 4.16), and hence easier to understand than the motive of each $\textup{BT}_{n}$. To transfer information from Disp to BT, a definition of the motive of a pro-system of algebraic stacks should be compatible with morphisms, at least those of pro-(locally of finite type).
In this article, we will investigate such situations more closely. To be more precise, we will study the category of rational motives on pro-systems of algebraic stacks $X\simeq\lim_{i}X_{i}$ defined via $\mathop{\textup{colim}}_{i}\textup{DM}(X_{i})$, and investigate its functorial properties.
@@
Motives on algebraic stacks: We can define $\textup{DM}(F)$, the DG-category of motives for any prestack $SF\colon\textup{(Ring)}\rightarrow\SS$ (here $S\SS$ denotes the $\infty$-category of Kan-complexes) via right Kan extension along the Yoneda embedding [RS20]. In this way, we can define $\textup{DM}(X)$ for any algebraic stack locally of finite type over a field $K$ for rational motives. Since rational motives satisfy étale descent, the $\infty$-category of rational motives on an algebraic stack $X$ with smooth atlas $\tilde{X}$ can be computed via the formula
${\textup{DM}(X)\coloneqq\lim(\textup{DM}(\tilde{X})}$${\textup{DM}(\tilde{X}\times_{X}\tilde{X})}$$\dots)$}
[RS20]. The problem with these definitions is that the computation of the motive of a limit $\lim_{i}X_{i}$ as above is not straightforward. To circumvent this problem, we will work with an alternative definition and show that in classical situations our definition agrees with the usual one.
Motives on pro-algebraic stacks: Let us fix a filtered $\infty$-category $\mathcal{I}$}.
@@
Let $K$ be a field and $x\colon{\mathcal{I}}\rightarrow\textup{PreStk}_{K}^{\operatorname{op}}$ be an inverse system of prestacks. Assume that each $X_{i}\coloneqq x(i)$ is representable by an algebraic stack locally of finite type over $K$ and let us denote by $X\coloneqq\lim x$. Then we call the tuple $(x,X)$ a pro-$\mathcal{I}$}-algebraic stack.
We call $(x,X)$ tame, if $x_{ij}$ is smooth for all $i\to j\in{\mathcal{I}}$.
Note that there is an obvious notion of morphisms of pro-$\mathcal{I}$}-algebraic stacks.
@@
Let $(x,X)$ be a pro-$\mathcal{I}$}-algebraic stack. Then we define the $\infty$-category of rational motives on $(x,X)$ to be
$\textup{DM}(x,X)\coloneqq\mathop{\textup{colim}}_{i\in{\mathcal{I}},x^{*}}\textup{DM}(X_{i}),$
where the colimit is computed in $\textup{DGCat}_{\textup{cont}}\coloneqq\mathrm{Mod}_{{\mathcal{D}}(\mathbb{Q})}(\Pr^{L}_{\mathrm{st}})$.
With this construction, we will obtain a full $6$-functor formalism on pro-$\mathcal{I}$}-algebraic stacks, which is not provided by the right Kan extended version. The caveat is that this $6$-functor formalism does not exist for all morphisms of pro-$\mathcal{I}$}-algebraic stacks. We have to make restrictions depending if $\mathcal{I}$$\simeq\mathbb{N}$, the poset $(\mathbb{N},\leq)$ viewed as an $\infty$-category, or not. We will explain this in the main text of this article.
To keep this introduction light, we will only state a summarized version of the existence of a (motivic) $6$-functor formalism and refer for a precise version to the main text.
@@
The $\infty$-category $\textup{DM}(x,X)$ is closed symmetric monoidal, compactly generated and the assignment $(x,X)\mapsto\textup{DM}(x,X)$ is contravariantly functorial in $(x,X)$.
Let $f\colon(x,X)\rightarrow(y,Y)$ be a morphisms of tame pro-$\mathbb{N}$-algebraic stacks that is levelwise locally of finite type. Then $f^{*}\coloneqq\textup{DM}(f)$ is part of a $6$-functor formalism if we assume that the exchange map $f_{n+1!}x_{n}^{!}\rightarrow y_{n}^{!}f_{n!}$ induced by the diagram is
$X_{n+1}$$X_{n}$$Y_{n+1}$$Y_{n}$$\scriptstyle{x_{n}}$$\scriptstyle{f_{n+1}}$$\scriptstyle{f_{n}}$$\scriptstyle{y_{n}}$
an equivalence for all $n\in\mathbb{N}$. Moreover, if $f_{n}^{*}y_{n*}\to x_{n*}f_{n+1}^{*}$ is an equivalence, then $f^{*}$ defines a pullback formalism with respect to smooth maps.
Further, this $6$-functor formalism satisfies homotopy invariance, T-stability, localization and purity in a suitable sense.
Finally, if we assume that each square above is cartesian, the same results hold for any locally of finite type morphism of pro-$\mathcal{I}$}-algebraic stacks for any filtered $\infty$-category without any assumption on tameness and the exchange map.
The assumption on the base change map is satisfied, for example, if $(x,X)$ and $(y,Y)$ are tame and square above is cartesian. Surprisingly, the condition on $f_{n+1!}x_{n}^{!}\rightarrow y_{n}^{!}f_{n!}$ only appears if the square above is not cartesian, since otherwise, we can use the !/!/*-base change equivalence. More importantly, we will see that it also holds for the structure map $(\tau,\textup{Disp})\to\operatorname{Spec}(\mathbb{F}_{p})$, (where $\tau_{n}\colon\textup{Disp}_{n+1}\to\textup{Disp}_{n}$ denotes the truncation map on displays), for which each square is not cartesian. In particular, we can define the motive of Disp via !!-pull/push of the unit. Moreover, we will see that in the special case of Disp, we have continuity
$\textup{DM}(\textup{Disp})\simeq\textup{DM}(\tau,\textup{Disp})$
(see Proposition 4.15). By Theorem 3, we obtain adjoint functors $p_{\sharp}\dashv p^{*}\dashv p_{*}$ and $p_{!}\dashv p^{!}$ with the expected compatibilities for the structure map $p\colon\textup{Disp}\to\operatorname{Spec}(\mathbb{F}_{p})$. These do not follow formally from the definition of $\textup{DM}(\textup{Disp})$ via Kan extension.
The definition via $\textup{DM}(\tau,\textup{Disp})$ allows us to compute $M(\textup{Disp})$ more explicitly.
Theorem 4 (4.16).
The pullback along the truncation map $\textup{Disp}\rightarrow\textup{Disp}_{1}$ induces a fully faithful embedding.
$\textup{DM}(\textup{Disp}_{1})\hookrightarrow\textup{DM}(\textup{Disp})$
In particular, we have an equivalence of motives
$M(\textup{Disp})\simeq M(\textup{Disp}_{1})$
inside $\textup{DM}(\mathbb{F}_{p})$ and moreover it is contained in the full stable cocomplete subcategory of $\textup{DM}(\mathbb{F}_{p})$ generated by Tate motives.
As a consequence, working with pro-algebraic stacks, we see that $H^{*,*}(BT,\mathbb{Q})\cong H^{*,*}(\textup{Disp}_{1},\mathbb{Q})$ (see Corollary 4.19).
Generalization: Since most arguments in the proof of Theorem 3 apply in a more general context, we will frequently work with compactly generated $6$-functor formalisms. Although our primary interest lies in rational motives, we will formulate the necessary statements as generally as possible. This broader perspective encompasses both compactly generated pullback formalisms [DG22] and $6$-functor formalisms in the sense of Liu–Zheng [LZ17]. The advantage of this generality is that the key computations remain valid in a wider range of settings, such as pro-systems of ind-schemes. The étale derived category of such systems, for instance, is considered by Bouthier, Kazhdan and Varshavsky [BKV22].
@@
We start by defining pro-$\mathcal{I}$}-algebraic stacks and listing some examples. We define the $\infty$-category of rational motives on pro-$\mathcal{I}$}-algebraic stacks.
Afterward, we prove, that on certain classes of morphisms, this definition yields a (motivic) $6$-functor formalism.
Lastly, we will have a closer look at motivic cohomology for pro-$\mathcal{I}$}-algebraic stacks and apply our results to the stack of displays and arbitrary Galois extensions.